### v1

# Bonsai: Gradient-free Graph Distillation for Node Classification

Abstract
Graph distillation has emerged as a promising avenue to enable scalable training of Gnns by compressing the training dataset while preserving essential graph characteristics. Our study uncovers significant shortcomings in current graph distillation techniques. First, the majority of the algorithms paradoxically require training on the full dataset to perform distillation. Second, due to their gradient-emulating approach, these methods require fresh distillation for any change in hyperparameters or Gnn architecture, limiting their flexibility and reusability. Finally, they fail to achieve substantial size reduction due to synthesizing fully-connected, edge-weighted graphs. To address these challenges, we present Bonsai, a novel graph distillation method empowered by the observation that computation trees form the fundamental processing units of message-passing Gnns. Bonsai distills datasets by encoding a careful selection of exemplar trees that maximize the representation of all computation trees in the training set. This unique approach imparts Bonsai as the first linear-time, model-agnostic graph distillation algorithm for node classification that outperforms existing baselines across $6$ real-world datasets on accuracy, while being $22$ times faster on average. Bonsai is grounded in rigorous mathematical guarantees on the adopted approximation strategies making it robust to Gnn architectures, datasets, and parameters.
1 Introduction and Related Works
Graph Neural Networks (Gnns) have shown remarkable success in various predictive tasks on graph data (Veličković et al., 2018; Kipf & Welling, 2017; Hamilton et al., 2017). However, real world graphs often contain millions of nodes and edges making the training pipeline slow and computationally demanding (Hu et al., 2021). This limitation hinders their adoption in resource-constrained environments (Miao et al., 2021) and applications dealing with massive datasets (Hu et al., 2021). Graph distillation (also called condensation) has emerged as a promising way to bypass this bottleneck (Jin et al., 2021; Liu et al., 2024a; b; Zheng et al., 2023). The objective in graph distillation is to synthesize a significantly smaller distilled data set that retains the essential information of the original data. By training Gnns on these distilled datasets, we can achieve comparable performance while reducing computational overhead and storage costs. This makes Gnns more accessible and practical for a wider range of applications, including those with limited computational resources and large-scale datasets. In this work, we study the problem of graph distillation for node classification.
1.1 Existing works and their Limitations
Table 1 presents existing graph distillation algorithms proposed for node classification and their characterization across various dimensions. We omit listing DosCond (Jin et al., 2022), Mirage (Gupta et al., 2024), and KiDD (Xu et al., 2023) in Table 1 since they are designed for graph classification, whereas we focus on node classification. The visual representation in Table 1 allows for quick identification of strengths and limitations of existing methods. We elaborate these limitations.
$\bullet$ Full Gnn training is a pre-requisite: The fundamental requirement of data distillation is that it should require less computational resources and time than training on the full dataset. However, this basic premise is violated by majority of the techniques (See Table 1) since they adopt a design where training the target Gnn on the full trainset is a prerequisite to distillation. These algorithms adopt a gradient-dependent optimization framework. They first train a Gnn on the full trainset and extract gradients of model parameters over epochs (Sfgc and Gcsr require training on the full dataset up to 200 and 100 times, respectively). The distillation process is formulated as an optimization problem to create a distilled dataset that replicates the gradient trajectory observed in the original training set. This need for full dataset training for distillation contradicts its fundamental premise, which we address in our work.
Table 1: Characterization of existing graph distillation algorithms. Cells shaded in Red indicate the presence of an undesirable property, while Green represents their absence.
Algorithm Requires training Gnn on full-dataset Generates fully-connected graph Model-specific distillation
GCond (Jin et al., 2021)
Sgdd (Yang et al., 2023)
Sfgc (Zheng et al., 2023)
Gdem (Liu et al., 2024a)
Gcsr (Liu et al., 2024b)
Bonsai
Table 2: We present the number of edges in the distilled graphs produced by GCond (Jin et al., 2021) and compare them to the full dataset. Red cells indicate cases where the distilled graph have more edges than the full dataset. The indicated node ratios are taken from the values used in GCond. Since all other baselines also produce fully-connected graphs, the numbers from this table translate to them as well.
Dataset Distilled Full Dataset
($r_{n}$ or node ratio) edge ratio #edges #edges
Cora (5.2%) 186% 19,600 10,556
Citeseer (3.6%) 158% 14,400 9,104
Pubmed (0.3%) 4% 3,600 88,648
Flickr (1%) 88% 795,664 899,756
Ogbn-arxiv (0.5%) 31% 715,716 2,315,598
Reddit (0.2%) 1% 216,225 23,213,838
$\bullet$ Distilling to a fully-connected graph: In a message-passing Gnn, the computation cost of each forward pass is $\mathcal{O}(|\mathcal{E}|)$, where $\mathcal{E}$ denotes the set of edges. Consequently, the computational effectiveness of graph distillation is primarily determined by the reduction in edge count between the original and distilled graphs, rather than node count alone. However, current graph distillation algorithms (see Table 1) quantify the condensation ratio based on the node count. Specifically, given a compression ratio $r_{n}$, it synthesizes a weighted, fully-connected dense adjacency matrix for the distilled graph of the size $\frac{|\mathcal{V}|}{r_{n}}\times\frac{|\mathcal{V}|}{r_{n}}$, where $\mathcal{V}$ denotes the node set. This design, which is inconsistent with the computation structure of Gnns, can lead to a distilled graph with minimal reduction in edge count and, in some cases, may even result in a graph with more edges than the original dataset (See Table 2).
$\bullet$ Model-specific distillation: Gradients of model weights are influenced by the specific Gnn architecture and hyper-parameters (e.g., number of layers, hidden dimensions, dropout rates). Consequently, any architectural change, such as switching from a Gcn (Kipf & Welling, 2017) to a Gat (Veličković et al., 2018), or hyper-parameter adjustments, necessitates a new round of distillation.
1.2 Contributions
To address the limitations outlined in Table 1, we present Bonsai  [^1] Inspired by the art of Bonsai, which transforms large trees into miniature forms while preserving their essence, our graph distillation algorithm gracefully prunes redundant computation trees, creating a distilled graph that is significantly smaller yet maintains comparable performance..
  • Gradient-free distillation: Instead of replicating the gradient trajectory, Bonsai emulates the distribution of input data processed by message-passing Gnns. By shifting the computation task to the pre-learning phase, Bonsai achieves independence from hyper-parameters and model architectures as long as it adheres to a message-passing Gnn framework like Gat (Veličković et al., 2018), Gcn (Kipf & Welling, 2017), GraphSage (Hamilton et al., 2017), Gin (Xu et al., 2019), etc. Moreover, this addresses a critical limitation of existing graph distillation algorithms that necessitates training on the entire training dataset.
  • Novel algorithm design: Bonsai is empowered by the observation that any message-passing Gnn decomposes a graph of $n$ nodes into $n$ rooted computation trees. Furthermore, topologically similar computation trees generate similar embeddings regardless of the Gnn being used (Xu et al., 2019; Togninalli et al., 2019). Bonsai exploits this observation to identify a small subset of diverse computation trees, called exemplars, that are located in dense regions and thereby representative of the full set. Hence, the induced subgraph spanned by the exemplars forms an effective distilled set.
  • Empirical evaluation: We perform rigorous benchmarking incorporating state-of-the-art graph distillation algorithms on $6$ real-world datasets containing up to hundreds of millions of edges. Our analysis establishes that Bonsai (1) achieves higher prediction accuracy, (2) produces $22$ times faster distillation times despite being CPU-bound in contrast to GPU-bound distillation of baselines, and (3) exhibits superior robustness to Gnn architectures and datasets.
2 Problem Formulation and Preliminaries
Definition 1 (Graph).
$\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$ denotes a graph over a finite, non-empty node set $\mathcal{V}$ and edge set =$\mathcal{E}=\{(u,v)\mid u,v\in\mathcal{\mathcal{V}}\}$. $\bm{X}\in\mathbb{R}^{\mid\mathcal{V}\mid\times\mid F\mid}$ denotes node attributes encoded using $F$-dimensional feature vectors. We denote the attributes of node $v$ as $\mathbf{x}_{v}$.
Two graphs are identical if they are isomorphic to each other.
Definition 2 (Graph Isomorphism).
Graph $\mathcal{G}_{1}$ is isomorphic to graph $\mathcal{G}_{2}$ if there exists a bijective mapping between their node sets that preserves both edge connectivity and node features. Specifically, $\mathcal{G}_{1}$ is isomorphic to $\mathcal{G}_{2}\iff\exists f:\mathcal{V}_{1}\rightarrow\mathcal{V}_{2}\text{ such that: }(1)\>f\text{ is a bijection, }(2)\>\mathbf{x}_{v}=\mathbf{x}_{f(v)},\text{ where }v\in\mathcal{V}_{1},f(v)\in\mathcal{V}_{2}\text{ and }(3)\>(u,v)\in\mathcal{E}_{1}\text{ if and only if }(f(u),f(v))\in\mathcal{E}_{2}$.
The problem of graph distillation for node classification is defined as follows.
Problem 1 (Graph Distillation).
Given train and validation graphs, $\mathcal{G}_{tr}$ and $\mathcal{G}_{val}$, respectively, and a memory budget $b$ in bytes, synthesize a graph $\mathcal{G}_{s}$ from $\mathcal{G}_{tr}$ within the budget, while minimizing the error gap between $\mathcal{G}_{s}$ and $\mathcal{G}_{tr}$ on the validation set, i.e., minimize $\left\{\left|\epsilon_{\mathcal{G}_{s}}-\epsilon_{\mathcal{G}_{tr}}\right|\right\}$. $\epsilon_{\mathcal{G}}$ represents the node classification error on the validation set when trained on graph $\mathcal{G}$.
2.1 Computation structure of Gnns
Gnns operate through an iterative process of information exchange between nodes. Let $\mathbf{x}_{v}\in\mathbb{R}^{\mid F\mid}$ represent the initial feature vector of node $v\in\mathcal{V}$. The propagation mechanism proceeds as follows:
[…]
Message creation: In layer $\ell$, collect and aggregate messages for each neighbor.
$\mathbf{m}_{v}^{\ell}(u)$ $=\text{{Message}}^{\ell}(\mathbf{h}_{u}^{\ell-1},\mathbf{h}_{v}^{\ell-1}),\quad\forall u\in\mathcal{N}_{v}=\{u\mid(u,v)\in\mathcal{E}\}$
$\overline{\mathbf{m}}_{v}^{\ell}$ $=\text{{Aggregate}}^{\ell}(\{\!\!\{\mathbf{m}_{v}^{\ell}(u):u\in\mathcal{N}_{v}\}\!\!\})$
[…]
A computation tree, also known as a receptive field, describes how information propagates through the graph during the neural network’s operation (Xu et al., 2019). Formally, it is defined as follows.
Figure 1: The figure depicts the construction of computation trees for nodes $v_{1}$ and $v_{11}$ in the sample graph, at depth $L=2$. Node colors indicate their labels. Despite being distant from each other in the graph and embedded in non-isomorphic $L$-hop neighborhoods, $v_{1}$ and $v_{11}$ have isomorphic computation trees.
Definition 3 (Computation Tree).
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$, a node $v\in\mathcal{V}$ and the number of layers $L$ in the Gnn, the computation tree $T^{L}_{v}$ rooted at $v$ is constructed as follows:
  • Enumerate all paths of length $L$ (including those with repeated vertices) starting from $v$.
  • Merge these paths into a tree structure where:
    1. 1.
      The root is always $v$, and
    2. 2.
      Two nodes from different paths are merged if: (i) They are at the same depth in their respective paths, and (ii) All their ancestors in the paths have already been merged.
Fig. 1 illustrates the idea of a computation tree with an example.
Properties of computation trees: We now highlight two key properties of computation trees, formally established in Xu et al. (2019), that form the core of our algorithm design. These properties hold regardless of the underlying message-passing Gnn (Gcn, Gat, GraphSage, Gin, etc.).
Property 1 (Sufficiency).
In an $L$-layered Gnn, the computation tree $\mathcal{T}_{v}^{L}$ is sufficient to compute node embedding $\mathbf{h}^{L}_{v}$, $\forall v\in\mathcal{V}$. Hence, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$, we may treat it as a (multi)set of computation trees $\mathbb{T}=\{\mathcal{T}_{v}^{L}\mid\forall v\in\mathcal{V}\}$. Clearly, $\left\lvert\mathbb{T}\right\rvert=\left\lvert\mathcal{V}\right\rvert$.
Property 2 (Equivalence).
If $\mathcal{T}_{v}^{L}$ is isomorphic to $\mathcal{T}_{u}^{L}$, then $\mathbf{h}_{v}^{L}=\mathbf{h}_{u}^{L}$. This follows from the property that the expressive power of a message-passing Gnn is upper bounded by the Weisfeiler-Lehman test (1-WL) (Xu et al., 2019). This implies that if the $L$-hop neighborhoods of two nodes are indistinguishable by 1-WL, then their representations will be the same. The 1-WL test cannot differentiate between isomorphic computation trees (Shervashidze et al., 2011).
Motivated with the above observations, we explore a relaxation of the Equivalence property: do nodes rooted at topologically similar computation trees generate similar node embeddings? To explore this hypothesis, we need a method to quantify the distance between computation trees. Given that the Equivalence property is derived from the 1-WL test, the Weisfeiler-Lehman kernel (Togninalli et al., 2019) presents itself as the natural choice for this distance metric.
Definition 4 (Weisfeiler-Lehman (WL) Kernel (Togninalli et al., 2019)).
Given graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$, WL-kernel constructs unsupervised embeddings over each node in the graph via a message-passing aggregation. Like in Gnns, the initial embedding $\mathbf{a}_{v}^{0}$ in layer 00 is initialized to $\mathbf{x}_{v}$. Subsequently, the embeddings in any layer $\ell$ is defined as:
$\mathbf{a}^{\ell}(v)=\frac{1}{2}\left(\mathbf{a}^{\ell-1}_{v}+\frac{1}{\deg(v)}\sum_{u\in\mathcal{N}(v)}w((v,u))\cdot\mathbf{a}^{\ell-1}_{u}\right)$ (1)
Figure 2: Pearson correlations between the $L_{2}$ distances of node pairs in the Gnn embedding space and the unsupervised embeddings derived from the WL-kernel, computed for pairs with a specific distance threshold in the WL-space ($x$-axis).
Here, $\mathcal{N}(v)$ denotes the neighbors of $v$, $\deg(v)=\left\lvert\mathcal{N}(v)\right\rvert$ and $w((v,u))=1$ for unweighted graphs.
By emulating the same message-passing framework employed by Gnns, the unsupervised embedding $\mathbf{a}_{v}^{L}$ jointly encapsulates the topology and node attributes within the computation tree $\mathcal{T}_{v}^{L}$. Hence, the distance between two computation trees $\mathcal{T}_{v}^{L}$ and $\mathcal{T}_{u}^{L}$ is defined to be:
$d\left(\mathcal{T}_{v}^{L},\mathcal{T}_{v}^{L}\right)=\left\|\mathbf{a}_{v}^{L}-\mathbf{a}_{u}^{L}\right\|_{2}$ (2)
While we use the $L_{2}$ norm, one may use other distance functions over vectors as well. The following corollary follows from the Eq. 2.
Corollary 1.
If $\mathcal{T}_{v}^{L}$ is isomorphic to $\mathcal{T}_{u}^{L}$, then $\mathbf{a}_{v}^{L}=\mathbf{a}_{u}^{L}$.
Hypothesis 1 (Relaxed Equivalence).
If $\mathcal{T}_{v}^{L}\sim\mathcal{T}_{u}^{L}$, then $\mathbf{a}_{v}^{L}\sim\mathbf{a}_{u}^{L}$, and, therefore $\mathbf{h}_{v}^{L}\sim\mathbf{h}_{u}^{L}$ irrespective of the specific message-passing architecture employed.
To test our hypothesis, in Fig. 2, we examine all node pairs within a specified distance threshold based on the Weisfeiler-Lehman (WL) kernel. We then analyze the correlation between their unsupervised WL-distance and the $L_{2}$ distance between their Gnn embeddings (Gcn with $2$ layers, trained with cross-entropy loss). The results reveal a compelling trend: as we decrease the distance threshold, we observe a strengthening correlation between WL-distance and Gnn embedding distance. This pattern strongly supports our hypothesis, indicating that computation trees that are proximate in the WL-space also exhibit proximity in the Gnn embedding space.
3 Bonsai: Proposed Methodology
Fig. 3 presents the pipeline of Bonsai, which is founded on the following logical progression:
  1. 1.
    Similar computation trees produce similar Gnn embeddings (Hypotheses 1).
  2. 2.
    Similar Gnn embeddings generate comparable outputs, resulting in similar impacts on the loss function and, consequently, on the gradients.
Figure 3: Pipeline of the proposed algorithm for graph distillation
Building on this reasoning, Bonsai aims to identify a set of $b$ exemplar computation trees that optimally represent the full training set [^2] We overload the notation for the budget by denoting the number of exemplar trees that fit within the input budget constraint (Recall Problem. 1) as $b$.. These exemplars are selected based on two critical criteria:
  • Representativeness: Each exemplar should be similar to a large number of other computation trees from the input training set, ensuring it captures common structural patterns. We quantify the representative power of a computation tree using the idea of reverse $k$ nearest neighbors (§ 3.1).
  • Diversity: The set of exemplars should be diverse, maximizing the coverage of different structural patterns present in the original graph (§ 3.2).
By optimizing for both representativeness and diversity, Bonsai ensures that the union of computation trees similar to the chosen exemplars encompasses a broad and representative spectrum of the original graph’s structural properties. The induced subgraph spanned by the nodes in these exemplar trees forms the initial distilled set. This initial version undergoes further refinement through an iterative process of edge sparsification and enrichment to reduce $O(|\mathcal{E}|)$ complexity of Gnns.
3.1 Quantifying Representativeness through Reverse $k$-NN
Utilizing Sufficiency (Property 1), we decompose the input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$ into a set of $\left\lvert\mathcal{V}\right\rvert$ computation trees $\mathbb{T}$ and then embed them into a feature space using WL-kernel (See Fig. 3). We now identify representative exemplars by analyzing this space.
The $k$ nearest neighbors of a computation tree $\mathcal{T}_{v}^{L}\in\mathbb{T}$, denoted as $k$-NN$(\mathcal{T}_{v}^{L})$, are its $k$ nearest computations trees from $\mathbb{T}$ in the WL-kernel space (Def. 4).
Definition 5 (Reverse $k$-NN and Representative Power).
[…]
$\Pi\left(\mathcal{T}_{v}^{L}\right)=\frac{\left\lvert\text{Rev-$k$-NN}(\mathcal{T}_{v}^{L})\right\rvert}{|\mathbb{T}|}$ (3)
If $\mathcal{T}_{v}^{L}\in k\text{-NN}(\mathcal{T}_{u}^{L})$ for a small $k$, it suggests that $\mathbf{h}_{v}^{L}\sim\mathbf{h}_{u}^{L}$ (Hypothesis 1). Consequently, a high $\Pi\left(\mathcal{T}_{v}^{L}\right)$ indicates that $\mathcal{T}_{v}^{L}$ frequently appears in the $k$-NN lists of many other trees and is therefore positioned in a dense region capturing shared characteristics across many trees. Hence, it serves as a strong candidate of being an exemplar.
3.1.1 Sampling for Scalable Computation of Reverse $k$-NN
Computing $k$-NN for each tree consumes $\mathcal{O}(n\log k)\approx\mathcal{O}(n)$ time (since $k\ll n$), where $n=\left\lvert\mathcal{V}\right\rvert$. Since $\left\lvert\mathbb{T}\right\rvert=n$, computing $k$-NN for all trees consumes $\mathcal{O}(n^{2})$ time. Hence, computing reverse $k$-NN for all trees consumes $\mathcal{O}(n^{2})$ time as well, which may be prohibitively expensive for graphs containing millions of nodes. To address this challenge, we employ a sampling technique that offers a provable approximation guarantee on accuracy. This is achieved as follows. Let us sample $z\ll n$ trees uniformly at random from $\mathbb{T}$, which we denote as $\mathbb{S}$. Now, we compute the $k$-NN of only trees in $\mathbb{S}$, which incurs $\mathcal{O}(zn)$ computation cost. We next approximate reverse $k$-NN of all trees in $\mathbb{T}$ based only on $\mathbb{S}$. Specifically, $\widetilde{\text{Rev-$k$-NN}}\left(\mathcal{T}_{v}^{L}\right)=\left\{\mathcal{T}_{u}^{L}\in\mathbb{S}\mid\mathcal{T}_{v}^{L}\in\text{$k$-NN}\left(\mathcal{T}_{u}^{L}\right)\right\}$. The approximated representative power is therefore:
$\widetilde{\Pi}\left(\mathcal{T}_{v}^{L}\right)=\frac{\left\lvert\widetilde{\text{Rev-$k$-NN}}(\mathcal{T}_{v}^{L})\right\rvert}{|\mathbb{S}|}$ (4)
The sample size $z$ balances the trade-off between accuracy and computational efficiency. By applying Chernoff bounds, we demonstrate in Lemma 1 that $z$ can be precisely determined to maintain the error within a specified error threshold $\theta$, with a confidence level of $1-\delta$.
Lemma 1.
Given a desirable error upper bound $\theta$ and a confidence interval of $\bm{1}-\delta$, if we sample at least $z=\frac{\ln{\left(\frac{2}{\delta}\right)}({2+\theta)}}{\theta^{2}}$ trees in $\mathbb{S}$, then for any $\mathcal{T}_{v}^{L}\in\mathbb{T}$:
$P(\left\lvert\widetilde{\Pi}(\mathcal{T}_{v}^{L})-\Pi\left(\mathcal{T}_{v}^{L}\right)\right\rvert\leq\theta)\geq 1-\delta$ (5)
Proof. For the formal proof, see App. A.1. Intuitively, if a tree participates in the $k$-NN of a large percentage of trees from the full dataset, then it will continue to do so in the sampled set as well. Lemma 1 establishes this formally resulting in the following positive implications.
  • The number of samples needed is independent of the size of $\mathbb{T}$.
  • Because $z$ grows logarithmically with $ln(\frac{2}{\delta})$, even a small sample size provides high confidence. Hence, the computation cost of Rev-$k$-NN reduces to $\mathcal{O}(n)$ since $z\ll n$.
3.2 Coverage Maximization
We aim to find the set of exemplar computation trees with the maximum representative power.
Definition 6 (The exemplars).
Let the representative power of a set of trees $\mathbb{A}$ be denoted by:
$\Pi\left(\mathbb{A}\right)=\left\lvert\bigcup_{\forall\mathcal{T}_{v}^{L}\in\mathbb{A}}\text{Rev-$k$-NN}\left(\mathcal{T}_{v}^{L}\right)\right\rvert\Biggm{/}{\lvert\mathbb{T}\rvert}$ (6)
Then, given the set of computation trees $\mathbb{T}$ and the budget $b$, we seek to identify the subset of computations trees, termed exemplars, $\mathbb{A}^{*}$, by maximizing the following objective:
$\mathbb{A}^{*}=\max_{\forall\mathbb{A}\subseteq\mathbb{T},\left\lvert\mathbb{A}\right\rvert=b}\Pi(\mathbb{A})$ (7)
For brevity of discourse, we proceed assuming the true reverse $k$-NN set.
Algorithm 1 The greedy approach
0:  $\mathcal{G}=(\mathcal{V},\mathcal{E},\bm{X})$, budget $b$
0:  solution set $\mathbb{A}$, $|\mathbb{A}|=b$
1:  $\mathbb{A}\leftarrow\emptyset$
2:  while $size(\mathbb{A})\leq b$ (within budget) do
3:     $\mathcal{T}_{v^{*}}^{L}\leftarrow\arg\max_{\mathcal{T}_{v}^{L}\in\mathbb{T}\backslash\mathbb{A}}\{\Pi(\mathbb{A}\cup\{\mathcal{T}_{v}^{L}\})-\Pi(\mathbb{A})\}$
4:     $\mathbb{A}\leftarrow\mathbb{A}\cup\{\mathcal{T}_{v^{*}}^{L}\}$
5:  Return $\mathbb{A}$
Theorem 1.
Maximizing the representative power of exemplars (Eq. 7) is NP-hard.
Proof. App. A.2 presents the formal proof by reducing to the Set Cover problem (Cormen et al., 2009). $\square$
Fortunately, Eq. 6 is monotone and submodular, which allows the greedy hill-climbing algorithm (Alg. 1) to approximate Eq. 7 within a provable error bound.
Theorem 2.
The exemplars, $\mathbb{A}_{greedy}$, selected by Alg. 1 provides an $1-1/e$ approximation, i.e., $\Pi(\mathbb{A}_{greedy})\geq\left(1-\frac{1}{e}\right)\Pi(\mathbb{A}^{*})$,
Proof. App. A.3 presents the proofs of monotonicity and submodularity of $\Pi(\mathbb{A})$. For monotonic and submodular functions, greedy selection provides $1-1/e$ approximation (Feige, 1998). $\square$.
We further enhance the efficiency of Alg. 1 by exploiting the property of monotonically decreasing marginal gains in submodular optimization (Leskovec et al., 2007).
Initial distilled graph: The initial distilled graph $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\bm{X}_{s})$ is formed by extracting the induced subgraph spanned by the exemplar computation trees, i.e., $\mathcal{V}_{s}=\left\{u\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}_{greedy},\>u\in\mathcal{N}_{L}(v)\right\}$, $\mathcal{E}_{s}=\{(u,v)\mid u\in\mathcal{V}_{s},v\in\mathcal{V}_{s}\}$ and $\bm{X}_{s}=\{\mathbf{x}_{v})\mid v\in\mathcal{V}_{s}\}$. Set $\mathcal{N}_{L}(v)$ contains nodes within $L$ hops from $v$.
3.3 Sparsification and Enrichment
Recall that the complexity of a Gnn forward pass is linear to the number of edges in the graph. In the final step, we seek to sparsify the graph induced by the exemplar trees. Furthermore, in the additional space created due to sparsfication, we include more exemplar trees, resulting in further magnification of the representative power. This process is formalized as follows.
Let $\mathcal{V}_{S}$ be the node set in the initial distilled graph, which can be partitioned into two disjoint subsets: $\mathcal{V}_{root}$ and $\mathcal{V}_{ego}$, where $\mathcal{V}_{root}=\{v\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}\}$, the root nodes, that serve as the root node of some exemplar in $\mathbb{A}_{greedy}$, and $\mathcal{V}_{ego}=\mathcal{V}_{S}\setminus\mathcal{V}_{root}$, the ego nodes, that are included because they fall within the $L$-hop neighborhood of a root node. We propose to employ Personalized PageRank (PPR) (Page et al., 1998) with teleportation only to root nodes to identify and prune ego nodes with minimal impact on the root node embeddings. Starting with timestamp $t=0$, we proceed as follows:
  1. 1.
    Compute the PPR distribution $\bm{\pi}_{t}$ for $t\geq 0$ defined by:
    $\bm{\pi}_{t}=(1-\beta)\mathbf{A}\bm{\pi}_{t-1}+\beta\mathbf{e}$ (8)
    where: $\bm{\pi}_{0}[i]=\frac{1}{|\mathcal{V}_{S}|}$ for all $i$, $\mathbf{e}[i]=\frac{1}{|\mathcal{V}_{root}|}$ if $v_{i}\in\mathcal{V}_{root}$, otherwise 00, $\mathbf{A}$ is the normalized adjacency matrix, $\beta$ is the teleportation probability.
  2. 2.
    Obtain the PPR vector $\bm{\pi}$ after convergence where $\bm{\pi}_{t}=\bm{\pi}_{t-1}$.
  3. 3.
    Sort the nodes in $\bm{\pi}$ by their PPR scores in descending order such that $\forall i,\>\bm{\pi}[i-1]\geq\bm{\pi}[i]$.
  4. 4.
    Define the knee point $i_{knee}$ as:
    $i_{knee}=\arg\max_{i}\{\bm{\pi}[i+1]+\bm{\pi}[i-1]-2\cdot\bm{\pi}[i]\}$ (9)
  5. 5.
    Prune all ego nodes with PPR scores below $\bm{\pi}[i_{knee}]$.
  6. 6.
    Use the freed-up space to include additional exemplar trees using Algorithm 1.
  7. 7.
    Repeat steps 1–6 iteratively until $i_{knee}\geq\theta$, where $\theta$ is a threshold setting the minimum number of nodes to be removed.
The rationale for this pruning process can be outlined as follows. The PPR distribution $\bm{\pi}_{t}$ iteratively updates based on the transition matrix $\mathbf{A}$ and teleportation vector $\mathbf{e}$. Given that $\bm{\pi}_{0}$ is initialized uniformly and $\mathbf{e}$ assigns higher probabilities to root nodes, $\bm{\pi}_{t}$ converges to a distribution where nodes more connected to root nodes have higher scores (See Fig. 6 in Appendix). After convergence, the vector $\bm{\pi}$ represents the steady-state probabilities of nodes in the graph. Sorting $\bm{\pi}$ by descending scores ensures that nodes with the highest influence on root nodes are prioritized. The knee point $i_{knee}$ is identified as the maximum curvature in the sorted PPR scores. This point is where the rate of change in scores shifts most significantly, indicating a transition between nodes of high and low influence. Nodes with PPR scores below $\bm{\pi}[i_{knee}]$ have minimal impact on root node embeddings and removing these nodes and edges incident on them effectively sparsifies the graph without substantial loss of information. Additional exemplar trees can then be incorporated into the available space using the greedy algorithm as given by Algorithm 1.
3.4 Properties of Bonsai
Complexity analysis: The computation compelxity of Bonsai is $\mathcal{O}(\lvert\mathcal{V}\rvert+\lvert\mathcal{E}\rvert)$. The detailed derivation is provided in App. A.4.
CPU-bound and Parallelizable: Bonsai does not involve learning any parameters through gradient descent or its variants. Hence, the entire process is CPU-bound. Furthermore, all steps are embarrassingly parallelizable leading distillation of datasets with hundreds of millions of edges within minutes.
Independence from model-architecture and hyper-parameters: Unlike majority of existing distillation algorithms that require knowledge of the Gnn architecture and all training hyper-parameters, Bonsai only requires approximate knowledge of the number of layers $L$. An approximate knowledge is sufficient since the influence of far-away neighbors gets squashed (Topping et al., 2022). For the same reason, Gnns are typically run with $L\leq 3$. Hence, we recommend setting $L\sim 2$.
4 Experiments
In this section, we benchmark Bonsai and establish:
  • Superior Accuracy: Bonsai consistently outperforms existing baselines in terms of accuracy across various compression factors, datasets, and Gnn architectures.
  • Enhanced Efficiency: On average, Bonsai is 350, 22, and 1500 times faster than the state-of-the-art baselines GCond, Gdem and Gcsr, respectively, in distilling datasets.
  • Increased Robustness: Unlike existing methods that require tuning distillation-specific hyper-parameters for each combination of Gnn architecture, dataset, and compression ratio, our approach surpasses all baselines using a single set of parameters across all scenarios.
Our implementation is available at https://anonymous.4open.science/r/bonsai.
4.1 Experimental Setup
The specifics of our experimental setup, including hardware and software environment, and hyper-parameters are detailed in App. B. For the baseline algorithms, we use the code shared by their respective authors. We conduct each experiment $5$ times and report the means and standard deviations. Across all datasets, we maintain a train-validation-test split ratio of $60:20:20$.
Table 3: Datasets.
Dataset # Nodes # Edges # Classes # Features
Cora (Kipf & Welling, 2017) 2,708 10,556 7 1,433
Citeseer (Kipf & Welling, 2017) 3,327 9,104 6 3,703
Pubmed (Kipf & Welling, 2017) 19,717 88,648 3 500
Flickr (Zeng et al., 2020) 89,250 899,756 7 500
Ogbn-arxiv (Hu et al., 2021) 169,343 2,315,598 40 128
Reddit (Hamilton et al., 2017) 232,965 23,213,838 41 602
Datasets: Table 3 lists the benchmark datasets used. The same datasets were also used in GCond, Gdem and Gcsr.
Baselines: Table 1 lists all the graph distillation algorithms for node classification. We compare Bonsai with GCond (Jin et al., 2021), Gdem (Liu et al., 2024a), and Gcsr (Liu et al., 2024b). We omit Sfgc (Zheng et al., 2023) and Sgdd (Yang et al., 2023) since both Gdem and Gcsr have been shown to outperform them. Among non-neural, baselines, we compare with selecting the induced subgraph spanned by a random selection of nodes, and Herding (Welling, 2009).
Metrics: Prediction quality is measured through Accuracy on the test set, i.e., the percentage of correct predictions. Compression ratio is quantified as $S_{r}=\frac{\text{size of distilled dataset in bytes}}{\text{size of full dataset in bytes}}$.
Table 4: Accuracy achieved by the various baselines on benchmark datasets across various compression ratios over byte consumption (denoted as $\mathbf{S_{r}(\%)}$), , on the Gcn architecture. The best and the second-best accuracies in each row are highlighted by dark and lighter shades of Green, respectively. OOT indicates the scenario where the algorithm failed to distill within 5 hours.
Dataset $\mathbf{S_{r}(\%)}$ Random Herding GCond GDEM GCSR Bonsai Full
0.5 $39.90{\scriptstyle\pm 1.46}$ $45.61{\scriptstyle\pm 0.01}$ $77.30{\scriptstyle\pm 0.30}$ $57.49{\scriptstyle\pm 6.87}$ $74.83{\scriptstyle\pm 0.75}$ $83.95{\scriptstyle\pm 0.39}$
Cora 1 $27.75{\scriptstyle\pm 2.05}$ $52.07{\scriptstyle\pm 0.00}$ $77.30{\scriptstyle\pm 0.31}$ $69.32{\scriptstyle\pm 4.71}$ $77.56{\scriptstyle\pm 0.74}$ $85.76{\scriptstyle\pm 0.24}$ $88.56{\scriptstyle\pm 0.18}$
3 $52.26{\scriptstyle\pm 1.69}$ $67.60{\scriptstyle\pm 0.00}$ $81.73{\scriptstyle\pm 0.48}$ $81.70{\scriptstyle\pm 3.10}$ $77.20{\scriptstyle\pm 0.48}$ $86.38{\scriptstyle\pm 0.22}$
0.5 $33.90{\scriptstyle\pm 2.16}$ $22.82{\scriptstyle\pm 0.00}$ $74.17{\scriptstyle\pm 0.68}$ $70.05{\scriptstyle\pm 2.40}$ $67.03{\scriptstyle\pm 0.61}$ $77.00{\scriptstyle\pm 0.15}$
CiteSeer 1 $44.90{\scriptstyle\pm 2.32}$ $49.10{\scriptstyle\pm 0.02}$ $77.62{\scriptstyle\pm 0.71}$ $72.48{\scriptstyle\pm 2.13}$ $74.77{\scriptstyle\pm 0.78}$ $77.03{\scriptstyle\pm 0.33}$ $78.53{\scriptstyle\pm 0.15}$
3 $44.50{\scriptstyle\pm 1.27}$ $67.69{\scriptstyle\pm 0.01}$ $77.02{\scriptstyle\pm 0.22}$ $76.20{\scriptstyle\pm 0.55}$ $77.27{\scriptstyle\pm 0.28}$ $75.89{\scriptstyle\pm 0.26}$
0.5 $62.58{\scriptstyle\pm 0.25}$ $78.29{\scriptstyle\pm 0.00}$ $80.63{\scriptstyle\pm 1.20}$ $80.72{\scriptstyle\pm 0.92}$ $79.43{\scriptstyle\pm 0.25}$ $87.27{\scriptstyle\pm 0.03}$
Pubmed 1 $79.19{\scriptstyle\pm 0.09}$ $78.59{\scriptstyle\pm 0.00}$ $79.92{\scriptstyle\pm 0.00}$ $80.80{\scriptstyle\pm 1.07}$ $79.11{\scriptstyle\pm 0.15}$ $87.08{\scriptstyle\pm 0.04}$ $87.22{\scriptstyle\pm 0.00}$
3 $82.50{\scriptstyle\pm 0.09}$ $78.09{\scriptstyle\pm 0.00}$ $77.00{\scriptstyle\pm 0.15}$ $81.07{\scriptstyle\pm 0.90}$ $79.94{\scriptstyle\pm 0.16}$ $87.64{\scriptstyle\pm 0.09}$
0.5 $44.78{\scriptstyle\pm 0.00}$ $47.98{\scriptstyle\pm 0.01}$ $44.06{\scriptstyle\pm 1.05}$ $46.25{\scriptstyle\pm 1.02}$ $46.41{\scriptstyle\pm 0.00}$ $48.73{\scriptstyle\pm 0.27}$
[…]
3 $46.56{\scriptstyle\pm 0.01}$ $46.54{\scriptstyle\pm 0.01}$ $46.04{\scriptstyle\pm 1.88}$ $47.35{\scriptstyle\pm 0.98}$ OOT $49.66{\scriptstyle\pm 0.27}$
0.5 $42.01{\scriptstyle\pm 0.01}$ $53.37{\scriptstyle\pm 0.00}$ $52.63{\scriptstyle\pm 0.63}$ $54.73{\scriptstyle\pm 0.66}$ OOT $58.49{\scriptstyle\pm 0.17}$
Ogbn-arxiv 1 $49.27{\scriptstyle\pm 0.64}$ $54.91{\scriptstyle\pm 0.00}$ $53.49{\scriptstyle\pm 0.63}$ $51.45{\scriptstyle\pm 1.14}$ OOT $58.35{\scriptstyle\pm 0.09}$ $68.97{\scriptstyle\pm 0.10}$
3 $51.11{\scriptstyle\pm 0.19}$ $57.28{\scriptstyle\pm 0.00}$ $53.01{\scriptstyle\pm 0.64}$ $53.37{\scriptstyle\pm 1.04}$ OOT $64.31{\scriptstyle\pm 0.06}$
0.5 $36.00{\scriptstyle\pm 4.09}$ OOT $38.94{\scriptstyle\pm 0.79}$ $90.51{\scriptstyle\pm 0.55}$ OOT $80.33{\scriptstyle\pm 0.46}$
Reddit 1 $38.55{\scriptstyle\pm 2.00}$ OOT $43.98{\scriptstyle\pm 0.35}$ $90.63{\scriptstyle\pm 0.84}$ OOT $85.65{\scriptstyle\pm 0.08}$ $92.14{\scriptstyle\pm 0.04}$
3 $44.97{\scriptstyle\pm 2.97}$ OOT $48.78{\scriptstyle\pm 0.83}$ $85.75{\scriptstyle\pm 0.80}$ OOT $88.90{\scriptstyle\pm 0.07}$
4.2 Prediction Accuracy
Table 4 presents the accuracies obtained by Bonsai and the baselines on Gcn. Several insights emerge from this experiment. Firstly, Bonsai achieves the best accuracy in 14 out of 18 scenarios, with substantial improvements ($\geq 5\%$) over the second-best performer in several cases. This demonstrates that the unsupervised gradient-agnostic approach adopted by Bonsai does not come at the cost of accuracy. Gdem emerges as the second best perfomer. Gcsr fails to complete in the two largest datasets of Ogbn-arxiv and Reddit since it performs training on the full dataset $100$ times to distill, which exceeds 5 hours.
Second, even at extreme compression ratio of $0.5\%$, it is possible to achieve reasonable accuracy levels. We highlight that past studies have denoted compression ratio in terms of number of nodes (recollect Table 2), which is misleading, since they generate fully connected weighted graphs, and thus real-compression in terms of bytes consumption, is lower.
4.3 Cross-architecture Generalization
Gdem and Bonsai are both model-agnostic, meaning their distilled datasets are independent of the downstream Gnn architecture used. In contrast, GCond and Gcsr produce architecture-specific distillation datasets, requiring separate training for each architecture. This distinction raises two important questions: First, how well do the distilled datasets generated by Gdem and Bonsai generalize across different architectures? Second, if we apply the distillation datasets produced by GCond and Gcsr for Gcn to other architectures, how significantly does it impact performance? We explore these questions in Table 5.
Consistent with the earlier trend, Bonsai continues to outperform all baselines. We further observe that the performance gap between Bonsai and the baselines is wider in Gat and Gin compared to Gcn (Table 4). For GCond and Gcsr, this wider gap is not surprising since they are trained on the gradients of Gcn, which are expected to differ from those of Gin and Gat. Gin uses a Sum-Pool aggregation, while Gat employs attention to effectively dampen unimportant edges, with a potentially different eigen spectrum - the property that Gdem attempts to preserve. The approach of Bonsai, on the other hand, aligns more directly with the computational structure of Gnns. It focuses on analyzing the input space of computation trees and aims to represent as much of this space as possible within the given budget. This approach of retaining essential information, rather than mimicking a pre-determined proxy function assumed to be important for Gnns, leads to superior results.
Table 5: Accuracies achieved by the various baselines on Gat and Gin at $1\%$ compression.
Dataset Gnn Random Herding GCond GDEM GCSR Bonsai Full
Gat $42.73{\scriptstyle\pm 1.03}$ $46.09{\scriptstyle\pm 0.86}$ $35.24{\scriptstyle\pm 0.00}$ $73.49{\scriptstyle\pm 2.64}$ $37.60{\scriptstyle\pm 1.34}$ $78.67{\scriptstyle\pm 0.89}$ $85.70{\scriptstyle\pm 0.09}$
Cora Gin $50.48{\scriptstyle\pm 0.85}$ $33.80{\scriptstyle\pm 2.42}$ $33.91{\scriptstyle\pm 1.23}$ $75.92{\scriptstyle\pm 4.24}$ $60.70{\scriptstyle\pm 4.44}$ $84.80{\scriptstyle\pm 0.41}$ $86.62{\scriptstyle\pm 0.28}$
Gat $46.19{\scriptstyle\pm 1.38}$ $52.07{\scriptstyle\pm 0.11}$ $21.47{\scriptstyle\pm 0.00}$ $23.87{\scriptstyle\pm 2.76}$ $21.50{\scriptstyle\pm 0.06}$ $69.43{\scriptstyle\pm 0.82}$ $77.48{\scriptstyle\pm 0.75}$
CiteSeer Gin $47.90{\scriptstyle\pm 0.65}$ $39.67{\scriptstyle\pm 0.82}$ $19.49{\scriptstyle\pm 1.09}$ $67.64{\scriptstyle\pm 4.45}$ $64.74{\scriptstyle\pm 1.88}$ $72.16{\scriptstyle\pm 0.60}$ $75.45{\scriptstyle\pm 0.23}$
Gat $78.85{\scriptstyle\pm 0.09}$ $76.64{\scriptstyle\pm 0.02}$ $41.55{\scriptstyle\pm 3.18}$ $80.75{\scriptstyle\pm 0.47}$ $40.47{\scriptstyle\pm 0.00}$ $85.88{\scriptstyle\pm 0.28}$ $86.33{\scriptstyle\pm 0.08}$
Pubmed Gin $78.43{\scriptstyle\pm 0.22}$ $62.22{\scriptstyle\pm 0.13}$ $32.84{\scriptstyle\pm 6.27}$ $78.72{\scriptstyle\pm 0.95}$ $33.75{\scriptstyle\pm 5.58}$ $85.57{\scriptstyle\pm 0.26}$ $84.66{\scriptstyle\pm 0.05}$
Gat $43.56{\scriptstyle\pm 1.06}$ $36.34{\scriptstyle\pm 1.14}$ $40.85{\scriptstyle\pm 1.08}$ $18.44{\scriptstyle\pm 9.42}$ OOT $45.62{\scriptstyle\pm 1.85}$ $51.42{\scriptstyle\pm 0.07}$
Flickr Gin $42.90{\scriptstyle\pm 0.76}$ $41.87{\scriptstyle\pm 4.52}$ $16.65{\scriptstyle\pm 6.55}$ $19.44{\scriptstyle\pm 9.68}$ OOT $44.90{\scriptstyle\pm 0.88}$ $45.37{\scriptstyle\pm 0.57}$
Figure 4: (a) Distillation times of various baselines on Gcn at a compression ratio of $0.5\%$, (b) Impact of the various components within Bonsai on its performance.
4.4 Distillation and Training Efficiency
Distillation efficiency: Fig. LABEL:fig:time presents the time consumed by Bonsai and other baselines. We note that while Bonsai is CPU-bound, all other algorithms are GPU-bound. Despite being CPU-bound, Bonsai, on average, is more than $22$-times faster than the closest baseline Gdem. Gdem has a quadratic time complexity due to being reliant on eigen decomposition. In contrast, Bonsai is linear to the number of nodes and edges in the graph. In addition, all three baselines require training on the full dataset. Gdem incurs a smaller cost by training an MLP solely on node attributes to predict node labels as a pre-processing step. However, GCond and Gcsr perform full Gnn training on the entire training set. This approach essentially negates the purpose of graph distillation, as it involves the very computational burden we aim to avoid. GCond is marginally faster than full training due to employing some Gcn-specific optimizations that outperform the native Gcn implementation available in PyG v2.3.1, which we use for full dataset training. Due to the same reason, Gcsr, although performing full-dataset training 100 times, does not incur a 100-fold increase in running time.
Training efficiency: Fig. 5 presents the training times on the three largest distilled datasets in Flickr, Ogbn-arxiv and Flickr. Results on other datasets are presented in Figs. 8, 9 and 10 in the App. Two key patterns emerge. Bonsai is faster across all architectures and datasets. Second, the gap between Bonsai and the baselines increase with increase in distilled dataset size. This is explained from the property that all baselines generate a fully connected weighted graph. Hence, the $L$-hop neighborhood of a node is significantly larger leading to higher computation and storage requirements. In contrast, through PPR, Bonsai sparsifies the graph by pruning the nodes with small impacts on embeddings of exemplars.
Component-wise analysis: We discuss the individual time consumption by each component of Bonsai in App. B.4.1.
Figure 5: Training times on the distilled datasets produced by Bonsai and the considered baselines on the three largest datasets at $0.5\%$ compression. Note that there is no bar for Gcsr in Ogbn-arxiv and Reddit since Gcsr fails to create a distilled dataset within 5 hours.
4.5 Ablation Study and Impact of Parameters
Fig. LABEL:fig:ablation presents an ablation study comparing the full Bonsai method against two variants: Bonsai-PPR, which omits the use of Personalized PageRank, and Bonsai-Rev-$k$-NN, which substitutes the Rev-$k$-NN-based coverage maximization with random exemplar selection. The results reveal that the Rev-$k$-NN component generally has a more substantial impact on performance, with the exception of the Reddit dataset where PPR marginally dominates.
This observation aligns with expectations, as random exemplar selection would fail to comprehensively represent the input space. The PPR component, while important, typically has a lesser impact since its primary function is to sparsify the initially distilled graph by eliminating message passing paths that minimally affect the embeddings of selected exemplars.
The Reddit dataset presents a unique case where PPR has a more significant impact. This is attributed to the dataset’s high average node degree of 100100~{}100100, in contrast to less than $14$ in other networks. Consequently, within the given budget, very few exemplars can be selected due to their substantially larger $L$-hop neighborhoods. In this scenario, PPR plays a crucial role by pruning these expansive neighborhoods, thereby creating space for additional exemplars and improving overall performance.
Impact of Parameters: We analyze the impact of sampling size and $k$ on Rev-$k$-NN in App. B.4.2.
5 Conclusions and Future Works
In this work, we have developed Bonsai, a novel graph distillation method that addresses critical limitations in existing approaches. By leveraging the fundamental role of computation trees in message-passing Gnns, Bonsai achieves superior performance in node classification tasks across multiple real-world datasets. Our method stands out as the first linear-time, model-agnostic graph distillation algorithm, offering significant improvements in both accuracy and computational efficiency. Bonsai’s unique approach of encoding exemplar trees that maximize representation of the full training set’s computation trees has proven to be highly effective. This strategy not only overcomes the paradoxical requirement of training on full datasets for distillation but also eliminates the need for repeated distillation when changing hyper-parameters or Gnn architectures. Furthermore, Bonsai achieves substantial size reduction without resorting to fully-connected, edge-weighted graphs, thereby reducing computational demands. In contrast to baselines, Bonsai is completely CPU-bound, resulting in a smaller greenhouse footprint along with reduced computation demands, making it a more environmentally friendly option. These features collectively position Bonsai as a significant advancement in the field of graph distillation, offering both performance benefits and sustainability advantages.
Limitations and Future Works: While existing research on graph distillation has primarily focused on node and graph classification tasks, Gnns have demonstrated their versatility across a broader spectrum of applications. In our future work, we aim to expand the scope of graph distillation by developing task-agnostic data distillation algorithms.
Reproducibility Statement
To support the reproducibility of our work, we provide several resources in the paper and its supplementary materials. The source code of our models and algorithms is available at https://anonymous.4open.science/r/bonsai, which also includes details of how to train the baseline models in different settings. All theoretical results and assumptions are detailed in the Appendix A, ensuring clarity in our claims. Full experimental settings are documented in Appendix B.
Appendix A Appendix
A.1 Proof. of Theorem 1.
Recall the notation that $\mathbb{S}\subseteq\mathbb{T}$ denotes the subset of trees from which the representative power of all trees in $\mathbb{T}$ is being estimated. Let $\left\lvert\mathbb{S}\right\rvert=z$.
Corresponding to each $\mathcal{T}_{u}^{L}\in\mathbb{S}$, let us define an indicator random variable $X^{v}_{u}$ denoting whether $\mathcal{T}_{v}^{L}\in k\text{-NN}(\mathcal{T}_{u}^{L})$. Therefore, we have:
$X^{v}=\sum_{\forall\mathcal{T}_{u}^{L}\in\mathbb{S}}X^{v}_{u}=z\widetilde{\Pi}\left(\mathcal{T}_{v}^{L}\right)$ (10)
Since each tree in $\mathbb{S}$ is sampled uniformly at random from $\mathbb{T}$, we have $E(X^{v})=z\Pi\left(\mathcal{T}_{v}^{L}\right)$. We also note that $X^{v}\sim Binomial(z,\Pi\left(\mathcal{T}_{v}^{L}\right))$. Given any $\epsilon\geq 0$, from Chernoff bounds (Motwani & Raghavan, 1995), we have:
Upper tail bound:
$P\left(X^{v}\geq(1+\epsilon)z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)$ $\leq\exp{\left(-\frac{\epsilon^{2}}{2+\epsilon}z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)}$ (11)
Lower tail bound:
$P\left(X^{v}\leq(1-\epsilon)z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)$ $\leq\exp{\left(-\frac{\epsilon^{2}}{2}z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)}$ (12)
Combining the upper and tail bounds, we obtain:
$P\left(\left\lvert X^{v}-z\Pi\left(\mathcal{T}_{v}^{L}\right)\right\rvert\geq\epsilon z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)$ $\leq 2\exp{\left(-\frac{\epsilon^{2}}{2+\epsilon}z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)}$ (13)
[…]
Plugging $\theta$ from Eq. 5 in Eq. 15, we obtain $\theta=\epsilon\Pi\left(\mathcal{T}_{v}^{L}\right)$. Hence,
$P\left(\left\lvert\widetilde{\Pi}(\mathcal{T}_{v}^{L})-\Pi\left(\mathcal{T}_{v}^{L}\right)\right\rvert\geq\theta\right)$ $\leq 2\exp{\left(-\frac{\left(\frac{\theta}{\Pi\left(\mathcal{T}_{v}^{L}\right)}\right)^{2}}{2+\frac{\theta}{\Pi\left(\mathcal{T}_{v}^{L}\right)}}z\Pi\left(\mathcal{T}_{v}^{L}\right)\right)}$ (16)
$\leq 2\exp{\left(-\frac{\theta^{2}}{2\Pi\left(\mathcal{T}_{v}^{L}\right)+\theta}z\right)}$ (17)
For any $\mathcal{T}_{v}^{L}\in\mathbb{T},\;\Pi\left(\mathcal{T}_{v}^{L}\right)\leq 1$. Thus,
$P\left(\left\lvert\widetilde{\Pi}(\mathcal{T}_{v}^{L})-\Pi\left(\mathcal{T}_{v}^{L}\right)\right\rvert\geq\theta\right)\leq 2\exp{\left(-\frac{\theta^{2}}{2+\theta}z\right)}$ (18)
Now, if we want a confidence interval of $1-\delta$ (as stated in Eq. 5), we have:
$\delta\geq 2\exp{\left(-\frac{\theta^{2}}{2+\theta}z\right)}$ (19)
[…]
Hence, if we sample at least $\frac{\ln{\left(\frac{2}{\delta}\right)}({2+\theta)}}{\theta^{2}}$ trees in $\mathbb{S}$, then for any $\mathcal{T}_{v}^{L}\in\mathbb{T}$:
$\widetilde{\Pi}(\mathcal{T}_{v}^{L})\in[\Pi\left(\mathcal{T}_{v}^{L}\right)-\theta,\Pi\left(\mathcal{T}_{v}^{L}\right)+\theta]\text{ with probability at least }1-\delta.\hskip 36.135pt\square$ (25)
A.2 NP-hardness: Proof of Theorem 1
Proof. To establish NP-hardness of our proposed problem we reduce it to the classical Set Cover problem (Cormen et al., 2009).
Definition 7 (Set Cover).
Given a budget $b$ and a collection of subsets $\mathcal{S}=\{S_{1},\cdots,S_{m}\}$ from a universe of items $U=\{u_{1},\cdots,u_{n}\}$, i.e., $\forall S_{i}\in\mathcal{S},\>S_{i}\subseteq U$, the Set Cover problem seeks to determine whether there exists a subset $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ such that $|\mathcal{S}^{\prime}|=b$ and it covers all items in the universe, i.e., $|\bigcup_{\forall S_{i}\in S^{\prime}}S_{i}|=U$.
We show that given any instance of a Set Cover problem $\langle\mathcal{S},U\rangle,b$, it can be mapped to the problem of maximizing the representative power (Eq. 7).
[…]
With this construction, we can state that there exists a Set Cover of size $b$ if and only if there exists a solution set $\mathbb{A}\subseteq\mathbb{T}$ such that $\Pi(\mathbb{A})=\frac{\lvert\mathbb{T}_{U}\rvert}{\lvert\mathbb{T}\rvert}$. $\square$
A.3 Proofs of Monotonicity, Submodularity
Theorem 3 (Monotonicity).
$\forall$ $\mathbb{A}^{\prime}\supseteq\mathbb{A},\>\Pi(\mathbb{A}^{\prime})-\Pi(\mathbb{A})\geq 0$, where $\mathbb{A}$ and $\mathbb{A}^{\prime}$ are sets of computation trees.
Proof.
Since the denominator in $\Pi(\mathbb{A}$) is constant, it suffices to prove that the numerator is monotonic. This reduces to establishing that:
$\left\lvert\bigcup_{\forall\mathcal{T}_{v}^{L}\in\mathbb{A}^{\prime}}\text{Rev-$k$-NN}\left(\mathcal{T}_{v}^{L}\right)\right\rvert\geq\left\lvert\bigcup_{\forall\mathcal{T}_{v}^{L}\in\mathbb{A}}\text{Rev-$k$-NN}\left(\mathcal{T}_{v}^{L}\right)\right\rvert$ (26)
This inequality holds true because the union operation is a monotonic function. As $\mathbb{A}^{\prime}$ is a superset of $\mathbb{A}$, the union over $\mathbb{A}^{\prime}$ will always include at least as many elements as the union over $\mathbb{A}$. ∎
Theorem 4 (Submodularity).
$\forall\mathbb{A}^{\prime}\supseteq\mathbb{A},\>\Pi(\mathbb{A}^{\prime}\cup\{\mathcal{T}\})-\Pi(\mathbb{A}^{\prime})\leq\Pi(\mathbb{A}\cup\{\mathcal{T}\})-\Pi(\mathbb{A})$, where $\mathbb{A}$ and $\mathbb{A}^{\prime}$ are sets of computation trees and $\mathcal{T}$ is a computation tree.
Proof by Contradiction. Let us assume
$\exists\mathbb{A}^{\prime}\supseteq\mathbb{A},\>\Pi(\mathbb{A}^{\prime}\cup\{\mathcal{T}\})-\Pi(\mathbb{A}^{\prime})>\Pi(\mathbb{A}\cup\{\mathcal{T}\})-\Pi(\mathbb{A})$ (27)
Eq. 27 implies:
[…]
$\implies$ $\mathbb{A}\supseteq\mathbb{A}^{\prime},\text{ which is a contradiction.\hskip 216.81pt\hfill}\square$. □
Figure 6: Distribution of PPR scores against node ranks.
A.4 Complexity Analysis
For this analysis, we assume $\lvert\mathcal{V}\rvert=n$ and $\lvert\mathcal{E}\rvert=m$. In sync with Fig. 3, the computation structure of Bonsai can be decomposed into four components.
  1. 1.
    Embed trees into WL-space: This operation requires a pass through each twice, and is identical to the message-pass structure of a Gnn. Hence, the computation cost is bounded to $\mathcal{O}(m)$. A crucial distinction from a Gnn is that it does not require back-propagation and hence is fully CPU-bound.
  2. 2.
    Rev-$k$-NN: As discussed in § 3.1.1, computing Rev-$k$-NN consumes $\mathcal{O}(n)$ time with sampling.
  3. 3.
    Coverage Maximization: Each iteration in Alg. 1, iterates over all nodes (trees) in the graph ($\mathbb{T}$) that have not yet been added to the set of exemplars. For each node (equivalently, tree rooted at this node), we compute its marginal gain (line 3 in Alg. 1), which is bounded by the sample size $z$, since the cardinality of a Rev-$k$-NN set is bounded by the number of trees sampled for Rev-$k$-NN computation ((§ 3.1.1)). Since $z\ll n$, the complexity per iteration is $\mathcal{O}(zn)\approx\mathcal{O}(n)$. The number of iterations is bounded by the size of the distillation budget, which we typically set to less than $3\%$ of the full dataset and hence has negligible impact on the complexity.
  4. 4.
    PPR: PPR can be computed in linear time (Lofgren et al., 2014). Since the size of the distilled set is at most the full graph, this bounds the cost of PPT to $\mathcal{O}(n)$.
Combining all these components, the cost of Bonsai is bounded by $\mathcal{O}(n+m)$.
Appendix B Experimental Setup
B.1 Hardware Configuration
All experiments were conducted on a high-performance computing system with the following specifications:
  • CPU: 96 logical cores
  • RAM: 512 GB
  • GPU: NVIDIA A100-PCIE-40GB
B.2 Software Configuration
The software environment for our experiments was configured as follows:
  • Operating System: Linux (Ubuntu 20.04.4 LTS (GNU/Linux 5.4.0-124-generic x86_64)) )
  • PyTorch Version: 1.13.1+cu117
  • CUDA Version: 11.7
  • PyTorch Geometric Version: 2.3.1
B.3 Additional Experimental Parameters
  • Number of layers in evaluation models: 2 (with Relu in between) for Gcn, Gat, and Gin. The Mlp used in Gin is a simple linear transform with a bias defined by the following equation $\mathbf{W}\mathbf{X}+\mathbf{b}$ where $\mathbf{X}$ is the input design matrix.
  • Value of $k$ in Rev-$k$-NN: 5
  • Hyper-parameters Baselines: We use the config files shared by the authors. We note that the benchmark datasets are common between our experiments and those used in the baselines.
B.4 Additional Experiments
Table 6: Component-wise analysis of Bonsai’s running times in seconds.
Dataset WL Rev-$k$-NN Greedy PPR Full
Cora 0.87 0.05 0.01 1.68 2.61
Citeseer 0.97 0.07 0.02 1.71 2.77
Pubmed 8.89 1.39 0.64 13.93 24.85
Reddit 682.80 293.15 275.52 193.99 1445.45
B.4.1 Running time analysis of Bonsai
Fig. 6 presents the time consumed by the various components within Bonsai. We observe that most of the time is spent in PPR computation, except in Reddit, where Rev-$k$-NN is a dominant contributor. This trend can be explained by the high density of Reddit, where the average degree is 100100~{}100100 in contrast to $<5$ in the other networks.
Figure 7: Impact of $k$ in Rev-$k$-NN and sampling size for Rev-$k$-NN approximation on the performance of Bonsai.
B.4.2 Impact of Parameters on Bonsai
Bonsai has two parameters, both related to the computation of Rev-$k$-NN. The value of $k$ in Rev-$k$-NN and the sample size to accurately approximate the representative power.
Fig. LABEL:fig:samplingacc and Fig. LABEL:fig:samplingtime present the impact of sample size on accuracy and time, respectively. As expected, the general trend shows that increasing the sample size leads to higher accuracy and longer running times. However, we note that the increase in accuracy is minor. This observation is consistent with Lemma 1, which states that a small sample size is sufficient to achieve a high-quality approximation.
Fig. LABEL:fig:kvarus presents the impact of $k$ on accuracy. While the performance is generally stable against $k$, the peak accuracy is typically achieved when $k$ is set to 5.
B.5 Additional Training Times
Figure 8: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 0.5% of Cora, Citeseer, Pubmed, and Reddit. Note that there is no distilled dataset for Reddit output by Gcsr due to OOT.
Figure 9: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 1% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no distilled dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT.
Figure 10: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 3% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no distilled dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT; and distilled datasets produced by GCond and Gdem for Reddit cannot train on standard 2-layered PyTorch Geometric Gat and Gin due to OOM.
Fig. 8 presents the training times on distilled datasets in Cora, Citeseer, Pubmed, and Reddit at 0.5% size. Figs. 9 and 10 presents the training times on distilled datasets in Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit at 1%, and 3% sizes. Bonsai is faster on all architectures. The gap between Bonsai and the baselines increase with increase in distilled dataset size. This is more pronounced in larger datasets like Flickr, Ogbn-arxiv, and Reddit because the number of edges is $O(\lvert V\rvert^{2})$ for the baselines and forward pass through Gnns is usually $O(\lvert E\lvert)$. Additionally, the distilled dataset created by GCond and Gdem for Reddit at 3% size cannot be used to train basic PyTorch Geometric models for Gat and Gin, resulting in out-of-memory (OOM) errors because the backward pass graph maintained by PyTorch is larger.